5x+3y=300 3x+5y=600 find the value of x,y
Answer by Alan3354(69443)

5x+3y=300
3x+5y=600
---------
Multipy 1 eqn by 5 and the other by 3
---
25x +15y = 1500 - 1st eqn times 5
9x + 15y = 1800 - 2nd eqn times 3
---------------------------------------------- Subtract